OpenStudy (anonymous):

solve for X using x^2+bx=-c x^2-x+5 so x^2+8x=-5 but than im stuck

OpenStudy (anonymous):

where did you get the 8 from?

OpenStudy (anonymous):

woops accidently took that from a different problem im working on

OpenStudy (anonymous):

problem is x^2-x+5=0

OpenStudy (anonymous):

i think i use the equation x^2+b/a x = -c/a

OpenStudy (anonymous):

use the quadratic formula to solve

OpenStudy (anonymous):

or do you have to do it the way you have it up there ^^^?

OpenStudy (anonymous):

up there

OpenStudy (anonymous):

and that is where im getting lost. Let me try quadratic really fast

OpenStudy (anonymous):

quadratic works. I think i got it now lol

OpenStudy (anonymous):

oh wait no it doesnt because i get 1-20 within the rational...

OpenStudy (anonymous):

x^2-x+5=0 x^2-x=-5 x^2=-5+x x=sqrt(-5+x)

OpenStudy (anonymous):

so this problem has no solution than

OpenStudy (anonymous):

because i get 1±√1-20/2

OpenStudy (anonymous):

and the square cant be negative right

OpenStudy (anonymous):

that's the solution. it's just not exact. and squares shouldn't be negative.

OpenStudy (anonymous):

Hmm... ok thanks